OpenStudy (anonymous):

Rewrite the equation below in standard form. 2/3x=-5/2y+2 a) 4x+15y=12 b) 2x+15/2y=6 c) 2x+5y=4 d) 2x+5y=6

OpenStudy (anonymous):

a.

OpenStudy (anonymous):

2/3x = -5/2y + 2 --- multiply by the LCD, which is 6 4x = -15y + 12 -- add 15y 4x + 15y = 12

OpenStudy (anonymous):

thank u guys

OpenStudy (anonymous):

just remember, when you have fractions, if you multiply by the LCD, it will get rid of the fractions